OpenStudy (anonymous):

The very rigorous medical program at a local university has a 10% drop out rate for each year. If the school admits 1,000 freshmen, how many diplomas will need to be ordered 4 years later?

OpenStudy (anonymous):

just find 10% of 1000, then subtract it and find 10% of what's left, do that again 2 more times and you've got your answer

OpenStudy (anonymous):

so approximately 656 diplomas will need to be ordered

OpenStudy (anonymous):

You can also see it as: 90% stay each year 90% = 0.9 Current year * 0.9 = next year You're doing it for four years, so you'll be multiplying it by 0.9 4 times. 1000 * 0.9 * 0.9 * 0.9 * 0.9 = 1000 * 0.9^4 approx = 656

OpenStudy (anonymous):

thanks guys ! can you help me out with another one?

OpenStudy (anonymous):

yh! bluepig's way is much easier and quicker

OpenStudy (anonymous):

post it